# Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene

### v1

@@ VI Additional results: many-body projection
We have described how our many-body projection algorithm takes into account interacting effects from the integrated-out states. Then a non-projection, this is, a projection onto all bands, must return the original tight-binding system, $\mathcal{H}=H$, with the only difference that in $\mathcal{H}$ the density is measured with respect to $\rho_{0\bm{r}}$ (Eqs. (S12),(S13)) and in $H$ with respect to the uniform $1/2$ background (Eq. (S2)). $\mathcal{H}_{0}$ will then absorb the discrepancies in the definitions of the interaction. However, these differences are essentially negligible, as we can see in Fig. S4. This is consistent with the approximate relation $\rho_{0\bm{r}}=\sum_{\bm{k},n<0}|\langle\bm{r}|\bm{k}n\rangle|^{2}\approx\frac{1}{2}\sum_{\bm{k}n}|\langle\bm{r}|\bm{k}n\rangle|^{2}=\frac{1}{2}$.
Figure S4: Band structure of $\mathcal{H}_{0}$ in the non-projection ($n_{B}=11908$) model for $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0,5$ eV, compared to the non-interacting (bare) bands. They are almost identical.
@@
Figure S5: (a) Bare tight-binding, symmetric state and $\mathcal{H}_{0}$ ($n_{B}=4,20$) bands for coupling strength $\epsilon_{r}=50$, $U=0.5$ eV. Due to the weak coupling, the renormalized bands are only slightly wider. (b-c) The components of $h(\bm{k})$ of the flat-band theory ($n_{B}=4$) with (b) $\epsilon_{r}^{-1}=0$, $U=0$ eV (bare) and (c) $\epsilon_{r}=50$, $U=0.5$ eV.
VII Additional results: correlated states
@@
Assuming preserved moiré translations, the self-consistent states can be described by the one-particle density matrix, $P(\bm{k})_{\alpha\beta}=\frac{1}{2}\Big{(}\mathbb{1}_{\alpha\beta}+Q(\bm{k})_{\alpha\beta}\Big{)}=\langle c^{\dagger}_{\bm{k}\alpha}c_{\bm{k}\beta}\rangle,$ with the properties $P(\bm{k})^{2}=P(\bm{k})=P(\bm{k})^{\dagger}$ and $\text{tr}(P(\bm{k}))=$ the number of occupied electrons at $\bm{k}$. The indices $\alpha,\beta=(\tau\sigma s)$ denote the valley-sublattice-spin flavor and delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the expectation value. For spin-singlets and 4 occupied electrons at each $\bm{k}$, $Q(\bm{k})$ can be decomposed into products of $\sigma$ and $\tau$ matrices with real coefficients, $Q(\bm{k})=\sum_{ij}\langle\sigma_{i}\tau_{j}(\bm{k})\rangle\sigma_{i}\tau_{j}$. Under a unitary transformation $\mathcal{U}_{n}(\varphi)=\text{exp}(i\varphi\mathcal{G}_{n}/2)$, with $\mathcal{G}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}Gc_{\bm{k}}$, the $Q(\bm{k})$ matrix transforms to $U^{*}_{n}(\varphi)Q(\bm{k})U^{T}_{n}(\varphi)$ with $U_{n}(\varphi)=\text{exp}(i\varphi G_{n}/2)$. We obtained self-consistent Hartree-Fock states in the flat-band ($n_{B}=4$) system on a $24\times 24$ momentum grid; the band structures were computed afterwards on a finer grid.
@@
In Fig. S6 we study the KIVC (initial seed $Q(\bm{k})=\sigma_{y}\tau_{y}$ on the self-consistency loop), OP ($Q(\bm{k})=\sigma_{z}$), VP ($Q(\bm{k})=\tau_{z}$) and TIVC ($Q(\bm{k})=\sigma_{x}\tau_{x}$) states. Notice that, for the initial seeds, $|\text{KIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{OP}\rangle$, $|\text{VP}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{TIVC}\rangle$, $|\text{KIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{flat}})|\text{VP}\rangle$ and $|\text{TIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{flat}})|\text{OP}\rangle$.
@@
In Fig. S6(a), KIVC and OP are insulating while VP and TIVC are gapless. This can be understood because KIVC and OP break $C_{2z}\mathcal{T}$ and open a gap at the $K$ points; while VP and TIVC preserve $C_{2z}\mathcal{T}$ and the Dirac cones do not gap out but split in energy. Given that the bandwidth of $\mathcal{H}_{0}$ is comparable to the interaction scale (the splitting at $K_{M}$), the system remains metallic.
@@
This can also be understood in a more algebraic way. The properties $[\sigma_{x}\tau_{x},\sigma_{x}]=[\tau_{z},\sigma_{x}]=[\sigma_{x}\tau_{x},\sigma_{y}\tau_{z}]=[\tau_{z},\sigma_{y}\tau_{z}]=0$ and $P(\bm{k})^{2}=P(\bm{k})$ disallow a smooth transition between the different orders ($\tau_{z}$, $\sigma_{x}\tau_{x}$ at the BZ boundaries and $\tau_{x}$, $\sigma_{y}\tau_{z}$ at the BZ center); It has to be discontinuous, i.e. through a band crossing. In the TIVC phase we find multiple additional order parameters like $\sigma_{y}$, $\sigma_{x}\tau_{z}$ and $\sigma_{y}\tau_{x}$ in a small region around the band crossing, and the transition is sharp, but smooth. The consequence is the band crossings become avoided crossing with small gaps. Moreover, the different $C_{2x}$ eigenvalues of the valence bands at $\Gamma_{M}$ and $M_{M}$ oblige a genuine crossing along the $\Gamma_{M}M_{M}$ line in the form of a Dirac cone [98]. On the other hand, the OP and KIVC states have $Q=\sigma_{z}$ and $Q=\sigma_{y}\tau_{y}$ respectively at the Brillouin zone boundaries. These matrices anticommute with $\sigma_{x}$ and $\sigma_{y}\tau_{z}$ and thus a continuous transition between the different order parameters is allowed; the self-consistent OP and KIVC states are insulators with a smooth $Q$ matrix (modulo the vortex at $\Gamma_{M}$ coming from the gauge choice).
@@
In Fig. S5(b) we plot the main order parameters of KIVC and OP. The chiral symmetry relating the two phases is evident, as well as in Fig. S5(a) in the band structures of OP and KIVC, and VP and TIVC.
@@
In Fig. S6(c) we consider the self consistent KIVC state under flat and chiral transformations, $\mathcal{U}_{\text{flat}}(\varphi)|\text{KIVC}\rangle=\text{exp}(i\varphi/2\ \mathcal{G}_{\text{flat}})|\text{KIVC}\rangle$ and $\mathcal{U}_{\text{chiral}}(\varphi)|\text{KIVC}\rangle=\text{exp}(i\varphi/2\ \mathcal{G}_{\text{chiral}})|\text{KIVC}\rangle$. We plot the expectation value of the interaction ${\mathcal{H}}_{\text{int}}$ and the total energy ${\mathcal{H}}$ for varying $\varphi$. This provides an alternative confirmation that $\mathcal{H}_{0}$ $(={\mathcal{H}}-{\mathcal{H}}_{\text{int}})$ induces a strong $\mathcal{G}_{\text{flat}}$ breaking, while ${\mathcal{H}}_{\text{int}}$ is approximately $U(4)\times U(4)$ invariant. On the other hand, the smaller amplitude of the ${\mathcal{H}}_{\text{int}}$ curves of the flat rotations was expected form the larger singular values of $C_{2z}\overline{P}$ (Fig S2(b) and Table S2).
@@
In Fig. S7 we show the band structure and order parameters of the NSM, QAH and SP phases. The Dirac cones in NSM are visible in all panels, and the vortex at $\Gamma_{M}$ coming from the gauge choice is visible in the phase plot. The SP phase is metallic for the same reason as the VP.
@@
Figure S6: (a) Band structures of the Kramers intervalley coherent (KIVC), orbital polarized (OP), time-reversal invariant intervalley coherent (TIVC) and valley polarized (VP) states, for coupling strengths $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV. (b) Order parameters of KIVC and OP. (c) Expectation values of the energies of the rotated states $\mathcal{U}_{\text{flat}}(\varphi)|\text{KIVC}\rangle$ and $\mathcal{U}_{\text{chiral}}(\varphi)|\text{KIVC}\rangle$. The approximate chiral symmetry is manifest in (a-c) and the broken flat symmetry in (a).
@@
Figure S7: (a) Band structure of the valley $K$ valence band and order parameter of the nematic semimetal (NSM). (b) Band structure and order parameters of the Quantum Anomalous Hall (QAH) state. (c) Band structure of the spin polarized state (SP).

### v3

@@ Appendix F Additional information: many-body projection
On another note, we have described above how our many-body projection algorithm takes into account interacting effects from the integrated-out states. Consequently, a non-projection (this is, a projection onto all bands) must return the original tight-binding system, $\mathcal{H}=H$, with the only difference that in $\mathcal{H}$ the density is measured with respect to $\rho_{0\bm{r}}$ (Eqs. (16),(17)) and in $H$ with respect to the uniform $1/2$ background (Eq. (6)). $\mathcal{H}_{0}$ will then absorb the discrepancies in the definitions of the interaction. However, these differences are essentially negligible, as we can see in Fig. F4(c). This is consistent with the approximate relation $\rho_{0\bm{r}}=\sum_{\bm{k},n<0}|\langle\bm{r}|\bm{k}n\rangle|^{2}\approx\frac{1}{2}\sum_{\bm{k}n}|\langle\bm{r}|\bm{k}n\rangle|^{2}=\frac{1}{2}$.
In Fig. F5(a) we show the non-interacting (bare), the symmetric-state and the $\mathcal{H}_{0}$ (with $n_{B}=4,20$) band structures, for $\epsilon_{r}=50$, $U=0.5$ eV. In Fig. F5(b) we show the non-vanishing coefficients of $h(\bm{k})$ (Eq. 30) for the non-interacting (bare) and $n_{B}=4$ system with $\epsilon_{r}=50$, $U=0.5$ eV.
Figure F4: (a) Band structures of the symmetric state for $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV, showing the $20$ central bands in blue. (b) Gaps between the central and remote bands on the electron and hole side, for the states of valley $K$. On the left column there are the $6\times 6$ discretized Brillouin zone and on the right, the $96\times 96$ zone showing the band touchings near the $\Gamma_{M}M_{M}$ or symmetry-related lines. (c) Band structures of $\mathcal{H}_{0}$ in the non-projected ($n_{B}=11908$) model, compared to the non-interacting (bare) bands. They are almost identical.
@@
Figure F5: (a) Bare tight-binding, symmetric state and $\mathcal{H}_{0}$ ($n_{B}=4,20$) bands for coupling strength $\epsilon_{r}=50$, $U=0.5$ eV. Due to the weak coupling, the renormalized bands are only slightly wider. (b-c) The components of $h(\bm{k})$ of the flat-band theory ($n_{B}=4$) with (b) $\epsilon_{r}^{-1}=0$, $U=0$ eV (bare) and (c) $\epsilon_{r}=50$, $U=0.5$ eV.
Appendix G Additional results: correlated states
@@
Assuming preserved moiré translations, the self-consistent states can be described by the one-particle density matrix, $P(\bm{k})_{\alpha\beta}=\frac{1}{2}\Big{(}\mathbb{1}_{\alpha\beta}+Q(\bm{k})_{\alpha\beta}\Big{)}=\langle c^{\dagger}_{\bm{k}\alpha}c_{\bm{k}\beta}\rangle,$ with the properties $P(\bm{k})^{2}=P(\bm{k})=P(\bm{k})^{\dagger}$ and $\text{tr}(P(\bm{k}))=$ the number of occupied electrons at $\bm{k}$. The indices $\alpha,\beta=(\tau\sigma s)$ denote the valley-sublattice-spin flavor and delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the expectation value. For spin-singlets and 4 occupied electrons at each $\bm{k}$, $Q(\bm{k})$ can be decomposed into products of $\sigma$ and $\tau$ matrices with real coefficients, $Q(\bm{k})=\sum_{ij}\langle\sigma_{i}\tau_{j}(\bm{k})\rangle\sigma_{i}\tau_{j}$. Under a unitary transformation $\mathcal{U}_{n}(\varphi)=\text{exp}(i\varphi\mathcal{G}_{n}/2)$, with $\mathcal{G}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}Gc_{\bm{k}}$, the $Q(\bm{k})$ matrix transforms to $U^{*}_{n}(\varphi)Q(\bm{k})U^{T}_{n}(\varphi)$ with $U_{n}(\varphi)=\text{exp}(i\varphi G_{n}/2)$.
We obtained self-consistent Hartree-Fock states in the flat-band ($n_{B}=4$) system on a $24\times 24$ momentum grid. The wave functions of the flat bands were obtained by solving $H_{\text{MF}}$ (Eq. 7) with the normal-state Fock matrix of the original $6\times 6$ grid, on the $24\times 24$ grid. The band structures were obtained on finer grids the using the same method with the self-consistent Fock matrix.
@@
In Fig. G6 we study the KIVC (initial seed $Q(\bm{k})=\sigma_{y}\tau_{y}$ on the self-consistency loop), OP ($Q(\bm{k})=\sigma_{z}$), VP ($Q(\bm{k})=\tau_{z}$) and TIVC ($Q(\bm{k})=\sigma_{x}\tau_{x}$) states. Notice that, for the initial seeds, $|\text{KIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{OP}\rangle$, $|\text{VP}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{TIVC}\rangle$, $|\text{KIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{flat}})|\text{VP}\rangle$ and $|\text{TIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{flat}})|\text{OP}\rangle$.
@@
In Fig. G6(a), KIVC and OP are insulating while VP and TIVC are gapless. This can be understood because KIVC and OP break $C_{2z}\mathcal{T}$ and open a gap at the $K$ points; while VP and TIVC preserve $C_{2z}\mathcal{T}$ and the Dirac cones do not gap out but split in energy. Given that the bandwidth of $\mathcal{H}_{0}$ is comparable to the interaction scale (the splitting at $K_{M}$), the system remains metallic.
@@
This can also be understood in a different way. The properties $[\sigma_{x}\tau_{x},\sigma_{x}]=[\tau_{z},\sigma_{x}]=[\sigma_{x}\tau_{x},\sigma_{y}\tau_{z}]=[\tau_{z},\sigma_{y}\tau_{z}]=0$ and $P(\bm{k})^{2}=P(\bm{k})$ disallow a smooth transition between the different orders ($\tau_{z}$, $\sigma_{x}\tau_{x}$ at the BZ boundaries and $\tau_{x}$, $\sigma_{y}\tau_{z}$ at the BZ center); this transition has to be discontinuous, i.e. through a band crossing. In the TIVC phase we find multiple additional order parameters like $\sigma_{y}$, $\sigma_{x}\tau_{z}$ and $\sigma_{y}\tau_{x}$ in a small region around the band crossing, and the transition is sharp, but smooth. The consequence is the band crossings become avoided crossing with small gaps. Moreover, the different $C_{2x}$ eigenvalues of the valence bands at $\Gamma_{M}$ and $M_{M}$ oblige a genuine crossing along the $\Gamma_{M}M_{M}$ line in the form of a Dirac cone [102]. On the other hand, the OP and KIVC states have $Q=\sigma_{z}$ and $Q=\sigma_{y}\tau_{y}$ respectively at the Brillouin zone boundaries. These matrices anticommute with $\sigma_{x}$ and $\sigma_{y}\tau_{z}$ and thus a continuous transition between the different order parameters is allowed; the self-consistent OP and KIVC states are insulators with a smooth $Q$ matrix (modulo the vortex at $\Gamma_{M}$ coming from the gauge choice).
@@
In Fig. F5(b) we plot the main order parameters of KIVC and OP. The chiral symmetry relating the two phases is evident, as well as in Fig. F5(a) in the band structures of OP and KIVC, and VP and TIVC.
@@
In Fig. G6(c) we consider the self consistent KIVC state under flat and chiral transformations, $\mathcal{U}_{\text{flat}}(\varphi)|\text{KIVC}\rangle=\text{exp}(i\varphi/2\ \mathcal{G}_{\text{flat}})|\text{KIVC}\rangle$ and $\mathcal{U}_{\text{chiral}}(\varphi)|\text{KIVC}\rangle=\text{exp}(i\varphi/2\ \mathcal{G}_{\text{chiral}})|\text{KIVC}\rangle$. We plot the expectation value of the interaction ${\mathcal{H}}_{\text{int}}$ and the total energy ${\mathcal{H}}$ for varying $\varphi$. This provides an alternative confirmation that $\mathcal{H}_{0}$ $(={\mathcal{H}}-{\mathcal{H}}_{\text{int}})$ induces a strong $\mathcal{G}_{\text{flat}}$ breaking, while ${\mathcal{H}}_{\text{int}}$ is approximately $U(4)\times U(4)$ invariant. On the other hand, the amplitude of the ${\mathcal{H}}_{\text{int}}$ curves of the flat rotations being smaller than those of the chiral rotations was expected form the singular values of $C_{2z}\overline{P}$ being larger than the eigenvalues of $\overline{\sigma}_{z}$ (Fig D2(b) and Table D2).
@@
In Fig. G7 we show the band structure and order parameters of the NSM, QAH and SP phases. The Dirac cones in NSM are visible in all panels, and the vortex at $\Gamma_{M}$ coming from the gauge choice is visible in the phase plot. The SP phase is metallic for the same reason as the VP.
@@
Figure G6: (a) Band structures of the Kramers intervalley coherent (KIVC), orbital polarized (OP), time-reversal invariant intervalley coherent (TIVC) and valley polarized (VP) states, for coupling strengths $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV. (b) Order parameters of KIVC and OP. (c) Expectation values of the energies of the rotated states $\mathcal{U}_{\text{flat}}(\varphi)|\text{KIVC}\rangle$ and $\mathcal{U}_{\text{chiral}}(\varphi)|\text{KIVC}\rangle$. The approximate chiral symmetry is manifest in (a-c) and the broken flat symmetry in (a).
@@
Figure G7: (a) Band structure of the valley $K$ valence band and order parameter of the nematic semimetal (NSM). (b) Band structure and order parameters of the Quantum Anomalous Hall (QAH) state. (c) Band structure of the spin polarized state (SP).